OpenStudy (anonymous):

Please, I need help right now or Im pretty much dead! :( Question: A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (zzr0ck3r):

man you go to a tough school

OpenStudy (zzr0ck3r):

death? dear lord....

OpenStudy (zzr0ck3r):

so which once will produce the sequence 4,6,8,10.....?

OpenStudy (zzr0ck3r):

you know it must be the 2nd or 4th because the first term is 4

OpenStudy (anonymous):

Wouldn't it be the fourth?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Can you show me though how it works?

OpenStudy (anonymous):

Because that was pretty much a random guess of randomness lol

OpenStudy (zzr0ck3r):

well look at option c f(1) = 4 f(2) = f(1)+ 2(2) = 4+4=8

OpenStudy (zzr0ck3r):

we awnted 6

OpenStudy (zzr0ck3r):

but option d f(1) = 4 f(2) = f(2-1) + 2 = 4+2 = 6

OpenStudy (zzr0ck3r):

f(3) = f(3-1) + 2 = f(2) + 2 = 6+ 2 = 8

OpenStudy (zzr0ck3r):

you see?

OpenStudy (anonymous):

Yes, I do. Thank you so much!